Given functions 𝑓(π‘₯) = 6 + √π‘₯ βˆ’ 4 and 𝑔(π‘₯) =

10

π‘₯βˆ’5

.Β 

Determine the domain and range of 𝑓(π‘₯) and 𝑔(π‘₯)


1
Expert's answer
2022-03-18T16:54:17-0400

Functions have the following form: "f(x)=6+\sqrt{x-4}," "g(x)=10(x-5)". The domain of the first function is "[4,+\infty)", since the value under the square root must be nonnegative. The domain of the second function is "\mathbb{R}", since it is a linear function. The range of the first function is "[6,+\infty)", since the square root is always nonnegative. The range of the second function is "\mathbb{R}".